OpenStudy (anonymous):

The numbers on two consecutively numbered gym lockers have a sum of 131. What are the locker numbers?

OpenStudy (anonymous):

so let\[x + (x+1) = 131\] \[2x + 1 = 131\] \[2x = 130\] \[x = 65\] \[x+1 = 65 + 1\]